# Movement-Prediction-Adjusted Naïve Forecast

### v7

@@ 2.2 Movement Term Weight
Using this approach, we can determine the optimal value of $\theta$ and examine, through the difference between the MSEs, whether there is an improvement over the original naïve forecast under this optimal setting.
Suppose that the in-sample set has $N$ time steps and that movement prediction $\hat{D}_{t}$ is provided for each time step. Within the in-sample set, we assume that the movement prediction yields $n_{\text{in-sample}}^{\text{correct}}$ correct and $n_{\text{in-sample}}^{\text{incorrect}}$ incorrect directions. We subsequently used Eq. (2) to present the actual value and expand the MSE expression for the original naïve forecast:
$MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-y_{t-1}\right)^{2}\right)$ (6)
When using Eq. (2) to represent the actual value and Eq. (4) to represent the predicted value, the MSE of the adjusted naïve forecast on the in-sample set can be expressed as:
$MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$ (7)
where $\theta_{\text{in-sample}}$ represents the parameter $\theta$ estimated from the in-sample set and $\bar{\epsilon}_{\text{in-sample}}$ denotes the mean of $|\epsilon_{t}|$ calculated over the in-sample set. Consequently, the difference between the two MSEs on the in-sample set can be derived via Eq. (6) and Eq. (7), respectively:
$\Delta\text{MSE}_{\text{in-sample}}$ $=\text{MSE}_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (8)
$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx n_{\text{in-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in-sample}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx n_{\text{in-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in-sample}}$ (9)
Therefore, Eq. (8) can be simplified as follows:
$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (10)
@@
where $ACC_{\text{in-sample}}$ is the directional accuracy of the movement prediction on the in-sample set, which is defined as:
To maximize the $\Delta MSE_{\text{in-sample}}$, we focus on the expression inside parentheses in the final form of Eq. (10), which can be regarded as a function of parameter $\theta_{\text{in-sample}}$:
$f(\theta_{\text{in-sample}})=4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}}$ (12)
By differentiating $f(\theta_{\text{in-sample}})$ and setting $f^{\prime}(\theta_{\text{in-sample}})=0$, the critical value of $\theta_{\text{in-sample}}$, denoted as $\theta_{\text{in-sample}}^{*}$, is given by:
@@
Because $f^{\prime\prime}(\theta_{\text{in-sample}})<0$, the function $f(\theta_{\text{in-sample}})$ is concave downward, and the critical point of $f(\theta_{\text{in-sample}})$ is the maximum. Therefore, $\theta_{\text{in-sample}}^{*}$ represents the optimal value of parameter $\theta$ in the in-sample set. In addition, Eq. (13) reveals the relationship between the optimal value of parameter $\theta$ and the directional accuracy of the movement prediction within a fixed sequence length. According to Eqs. (10) and (13), the improvement over the original naïve forecast on the in-sample set, achieved by the adjusted naïve forecast with optimal parameters, is expressed as follows:
$\max\Delta MSE_{\text{in-sample}}$ $=\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$ (14)
According to Eqs. (13) and (14), as the directional accuracy increases from 0.5, the value of $f(\theta_{\text{in-sample}})$ increases from 0, and the improvement over the original naïve forecast correspondingly increases from 0. Because this improvement is a nonnegative value, the adjusted naïve forecast is at least as good as the original naïve forecast on the in-sample set:
@@
2.3 Forecast on the Out-of-Sample Set
If we use $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ as the estimates of the actual optimal values of $\theta$ and $\bar{\epsilon}$ on the out-of-sample set and suppose that the movement prediction on the out-of-sample set with $M$ time steps yields $m_{\text{out-of-sample}}^{\text{correct}}$ correct directions and $m_{\text{out-of-sample}}^{\text{incorrect}}$ incorrect directions; then, the difference between the two MSEs on the out-of-sample set can be expressed as follows:
$\Delta\text{MSE}_{\text{out-of-sample}}$ $=\text{MSE}_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (16)
Let $\bar{\epsilon}_{\text{out-of-sample}}$ denote the actual mean absolute increment on the out-of-sample set. According to the LLN, when both $m_{\text{out-of-sample}}^{\text{correct}}$ and $m_{\text{out-of-sample}}^{\text{incorrect}}$ are sufficiently large, the following approximations hold:
$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx m_{\text{out-of-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out-of-sample}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out-of-sample}}$ (17)
Therefore, Eq. (16) can further be expressed as:
@@ 2.3 Forecast on the Out-of-Sample Set
where $ACC_{\text{out-of-sample}}$ is the directional accuracy of the movement prediction on the out-of-sample set, which is defined as:
$ACC_{\text{out-of-sample}}=\frac{m_{\text{out-of-sample}}^{\text{correct}}}{m_{\text{out-of-sample}}^{\text{incorrect}}+m_{\text{out-of-sample}}^{\text{correct}}},$ (19)
and $\theta_{\text{out-of-sample}}^{*}$ is the actual optimal value of $\theta$ for the out-of-sample set, which is derived according to Eq. (13):
$\theta_{\text{out-of-sample}}^{*}=2\cdot ACC_{\text{out-of-sample}}-1$ (20)
To ensure the effectiveness of the adjusted naïve forecast, the $\Delta MSE_{\text{out-of-sample}}$ should be no less than zero. Therefore, according to Eq. (18), the following inequality must be satisfied:
@@
Given that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are specific estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, respectively, inequality (21) can be generalized as:
@@
where $\hat{\theta}_{\text{out-of-sample}}$ represents the estimate of $\theta_{\text{out-of-sample}}^{*}$ and $\hat{\bar{\epsilon}}_{\text{out-of-sample}}$ represents the estimate of $\bar{\epsilon}_{\text{out-of-sample}}$. Since the values of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ cannot be determined in advance, whether the inequality (22) can be satisfied is unknown, even when the optimal parameters derived from the in-sample set are used. Therefore, inequality (22) reveals the source of the uncertainty in the effectiveness of the adjusted naïve forecast, which depends on whether the product of the actual parameters exceeds half the product of the estimated parameters.
We can regard $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ as two variables, each of which follows a distribution with a nonnegative range. Inequality (22) represents an event whose probability of occurrence depends on the estimated values of the two variables. When both estimates are set to small values near zero, the probability that inequality (22) is satisfied approaches one. However, in such a setting, the improvement over the original naïve forecast by the adjusted naïve forecast on the out-of-sample set also diminishes. Therefore, inequality (22) also reflects a trade-off between maximizing the probability that the adjusted naïve forecast outperforms the original forecast and achieving a meaningful level of improvement.
In addition to using $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ as the estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, which may result in overfitting, a more conservative approach can be used to set the estimated values of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$:
@@
    Within each window, the directional accuracy is calculated and then translated into the corresponding value of $\theta$ via Eq. (13). As the window moves, a collection of $\theta$ values is generated, denoted as $\Theta_{\text{set}}$, whose bounds can be regarded as the estimated range of $\theta_{\text{out-of-sample}}^{*}$.
@@
    Similarly, a sliding window of the same length is applied to calculate a set of $\bar{\epsilon}$ values from the in-sample set, denoted as $\boldsymbol{\bar{\epsilon}}_{\text{set}}$, providing an estimated value range of $\bar{\epsilon}_{\text{out-of-sample}}$.
@@ 3 Simulation
This section presents a simulation study using four synthetic time series to evaluate whether the adjusted naïve forecast outperforms the original naïve forecast under different directional accuracy levels; therefore, movement prediction is assumed to have been provided. To statistically test the performance of the proposed method on an out-of-sample set, multiple movement predictions were randomly generated for each directional accuracy level.
@@ 3.1 Datasets and Setup
The dataset includes four synthetic time series, each of which represents a symmetric random walk with a certain type of variance, as summarized in Table 1. Each time series, consisting of 2,500 steps, used a fixed random seed (seed = 1) in the Python environment, as shown in Figure 2. To maintain all values within the positive range, an initial positive offset of 10000 units was applied to each series. The last 500 data points were selected as the out-of-sample set, and the remaining data points were used as the in-sample set. In addition, Figure 3 shows the first-order differences of each time series to visually demonstrate the properties of the increments, and Figure 4 shows the histograms of the first-order differences for all time series, highlighting the symmetric nature of the time series.
Table 1: Synthetic time series summary
@@
For each time series, the adjusted naïve forecast on the out-of-sample set was generated via Algorithm 1, which uses the values of $\theta$ determined by different levels of predefined directional accuracies. The performance of each adjusted naïve forecast result was evaluated using four metrics: the RMSE, MAE, MAPE, and sMAPE. The RMSE, derived from the square root of the MSE, emphasizes larger errors, which makes it suitable for identifying significant deviations in forecasts. The MAE, on the other hand, provides an interpretable measure of the average error without overpenalizing outliers. The MAPE and sMAPE normalize the errors relative to the actual values, enabling comparisons between datasets with different scales. Together, these metrics capture both absolute and relative error characteristics, ensuring a robust and balanced evaluation. The formulas for these metrics are as follows:
@@
To statistically test the performance of the adjusted naïve forecast, for each synthetic time series at a specified directional accuracy level, the process to generate the adjusted naïve forecast was repeated 100 times. Each repetition used a unique movement prediction sequence at the same directional accuracy level, which was generated by randomly replacing a subset of actual movements in the out-of-sample set with incorrect movements. The directional accuracy level varies in two ways: initially, it changes from 0.50 to 0.56 in increments of 0.01, and then it changes from 0.55 to 1.00 in larger increments of 0.05. This approach allows for finer granularity in the lower accuracy range, where small changes in the directional accuracy can have a significant effect on the performance of the method, while using larger steps at higher accuracy levels balances the computational efficiency. Consequently, for each synthetic time series, 100 independent values were calculated for each metric at each accuracy level.
@@ 3.2 Results and Analysis
Because the directional accuracy for the out-of-sample set is predefined in each simulation process, only the value of $\bar{\epsilon}_{\text{out-of-sample}}$ needs to be estimated. Figure 5 illustrates the distribution of the values in the $\boldsymbol{\bar{\epsilon}}_{\text{set}}$, generated on the basis of sliding windows across the first-order differences of the in-sample set. The vertical reference line in Figure 5 represents the actual $\bar{\epsilon}_{\text{out-of-sample}}$ value of the out-of-sample set. In the four cases, the minimum value of $\boldsymbol{\bar{\epsilon}}_{\text{set}}$ is generally an underestimate of $\bar{\epsilon}_{\text{out-of-sample}}$, except for the second case, where the synthetic time series is a random walk with a decreasing trend of linear variance. However, the estimated value remains close to the true value. Under the assumption that directional accuracy is consistent across the in-sample and out-of-sample sets, this estimation ensures that inequality (22) is satisfied.
@@
Furthermore, comparisons of the two types of naïve forecasts based on the four evaluation metrics at low accuracy levels are illustrated in Figures 6, 7, 8, and 9. Each box plot in each figure represents the distribution of 100 independent metric values at a specific directional accuracy level, with the metric value of the original naïve forecast shown as a reference line. According to Figure 6, at low accuracy levels, such as 0.51, the RMSE of the adjusted naïve forecast is comparable to or slightly better than that of the original naïve forecast, as shown by the proximity of the median of each box plot to the reference line. As the accuracy increased beyond 0.52, most of the RMSE values fell below the baseline.
@@
Because the adjusted naïve forecast consistently outperforms the original naïve forecast at low accuracy levels in terms of the MAE, MAPE, and sMAPE, the Wilcoxon signed rank test was conducted to specifically evaluate whether the median differences between the RMSE values of the adjusted naïve forecast and those of the original naïve forecast significantly deviated from zero for accuracy levels ranging from 0.50–0.56 (Li and Johnson,, 2014). As summarized in Table 2, at a directional accuracy level of 0.51, the adjusted naïve forecast shows mixed results with respect to statistical significance. Although significant improvements were observed over the original naïve forecast for the Type 4 series (randomly varying variance, p = 0.002), the results for the Type 1 (constant variance, p = 0.915), Type 2 (linear variance trend, p = 0.781), and Type 3 (cyclic variance, p = 0.812) series were not statistically significant, indicating that an accuracy level of 0.51 may not consistently produce meaningful improvements across all types of time series. As the accuracy increased beyond 0.51, significant reductions in the RMSE (p < 0.001) were observed for all types of time series, thereby demonstrating the reliability and effectiveness of the adjusted naïve forecast at higher accuracy levels.
@@
According to Figures 10, 11, 12, and 13, at higher accuracy levels ranging from 0.55–1.00, the adjusted naïve forecast consistently outperforms the original naïve forecast in terms of the four metrics. In particular, the improvement over the original naïve forecast becomes more pronounced as the directional accuracy improves. This finding demonstrates that highly accurate movement predictions can be effectively translated into improved point forecasts in the context of forecasting random walk-like time series. However, the improvement over the original naïve forecast is limited, because the proposed method allows a random walk-like series to be predicted to a certain extent, but the predicted value cannot be infinitely close to the actual value.
@@
In general, when the directional accuracy was 0.5, the adjusted naïve forecast did not provide an improvement over the original naïve forecast. However, as the directional accuracy increases, the performance of the adjusted naïve forecast improves. This result is consistent with the theoretical analysis of the proposed method, which highlights its robustness and potential applicability to a wide range of time series that exhibit symmetric random walk characteristics, including those with homoscedasticity as well as high-volatility scenarios.

### v8

@@ 2.2 Parameter Optimization Using In-Sample Data
Suppose the in-sample set has $N$ time steps, and movement predictions yield $n_{\text{in-sample}}^{\text{correct}}$ correct and $n_{\text{in-sample}}^{\text{incorrect}}$ incorrect directions; then, the directional accuracy of the movement prediction on the in-sample set can be expressed as follows:
Accordingly, the MSE difference between two types of naïve forecasts is as follows:
$\Delta MSE_{\text{in-sample}}$ $=\text{MSE}_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (7)
@@
where $\theta_{\text{in-sample}}$ represents the parameter $\theta$ in the adjusted naïve forecast for the in-sample set and $\bar{\epsilon}_{\text{in-sample}}$ denotes the mean of $|\epsilon_{t}|$ calculated over the in-sample set. A detailed algebraic derivation of the $\Delta MSE_{\text{in-sample}}$ is provided in Appendix A.
According to Eq. (7), the optimization reduces to maximizing the following function concerning parameter $\theta_{\text{in-sample}}$:
$f(\theta_{\text{in-sample}})=4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}},$ (8)
Differentiating Eq. (8) and setting the derivative to zero yields the optimal solution:
Proposition 1 (Optimal parameter solution). The optimal parameter value $\theta_{\text{in-sample}}^{*}$ that maximizes the MSE difference between the adjusted naïve forecast and the original naïve forecast on the in-sample set is given by:
@@
This result explicitly reveals that the optimal value of the parameter $\theta$ depends linearly on the directional accuracy of the movement predictions for a sequence with a fixed length. Since the second derivative of $f(\theta_{\text{in-sample}})$ is negative, this optimal solution is a maximum.
Substituting Eq. (9) back into Eq. (6) yields:
Proposition 2 (Performance guarantee). Under the optimal parameter setting using $\theta_{\text{in-sample}}^{*}$, the adjusted naïve forecast achieves a nonnegative improvement over the original naïve forecast in terms of the MSE on the in-sample set:
$\max\Delta MSE_{\text{in-sample}}=\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\geq 0$ (10)
Proposition 2 confirms that as the directional accuracy increases beyond 0.5, the value of $\theta_{\text{in-sample}}^{*}$ increases from 0, leading to progressively greater improvements over the original naïve forecast. At a directional accuracy of 1, the maximum achievable improvement is $\bar{\epsilon}_{\text{in-sample}}^{2}$. Thus, with the optimal parameter setting, the adjusted naïve forecast matches or surpasses the performance of the original naïve forecast within the in-sample dataset:
@@
This result underlines the theoretical robustness of the proposed adjustment methodology.
2.3 Uncertainty Analysis and Practical Generation of Out-of-Sample Forecasts
We now examine the performance of the adjusted naïve forecast via out-of-sample data. Suppose that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are used as the parameters of the adjusted naïve forecast on the out-of-sample set, which contains $M$ time steps. Suppose that the movement predictions on this set yield $m_{\text{out-of-sample}}^{\text{correct}}$ correct and $m_{\text{out-of-sample}}^{\text{incorrect}}$ incorrect predictions; then, the directional accuracy of the movement prediction on the out-of-sample set can be expressed as:
$ACC_{\text{out-of-sample}}=\frac{m_{\text{out-of-sample}}^{\text{correct}}}{m_{\text{out-of-sample}}^{\text{incorrect}}+m_{\text{out-of-sample}}^{\text{correct}}}$ (12)
Consequently, the MSE difference between two types of naïve forecasts on the out-of-sample is expressed as:
@@ 2.3 Uncertainty Analysis and Practical Generation of Out-of-Sample Forecasts
where $\bar{\epsilon}_{\text{out-of-sample}}$ denotes the actual mean absolute increment on the out-of-sample set and $\theta_{\text{out-of-sample}}^{*}$ denotes the actual optimal value of $\theta$ for the out-of-sample set, which is derived according to Proposition 1:
$\theta_{\text{out-of-sample}}^{*}=2\cdot ACC_{\text{out-of-sample}}-1.$ (14)
A detailed algebraic expansions of the $\Delta MSE_{\text{out-of-sample}}$ is provided in Appendix B. For the adjusted naïve forecast to outperform the original naïve forecast, the $\Delta MSE_{\text{out-of-sample}}$ should be no less than zero. Therefore, according to Eq. (13), the following inequality must hold:
@@
Given that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are specific estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, re2spectively, inequality (15) can be generalized as:
Proposition 3 (Out-of-sample performance criterion). The necessary condition for the adjusted naïve forecast to achieve an improvement over the original naïve forecast in terms of the MSE on the out-of-sample set is as follows:
@@
where $\hat{\theta}_{\text{out-of-sample}}$ represents the estimate of $\theta_{\text{out-of-sample}}^{*}$ and $\hat{\bar{\epsilon}}_{\text{out-of-sample}}$ represents the estimate of $\bar{\epsilon}_{\text{out-of-sample}}$.
Proposition 3 establishes a clear boundary criterion when an adjusted naïve forecast can outperform the original naïve forecast, specifically depending on whether the product of the actual optimal parameter values exceeds half the product of the estimated optimal parameter values. We can treat $\theta_{\text{out-of-sample}}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ as random variables, each following a distribution defined over nonnegative values. Accordingly, inequality (16) represents a probabilistic event, the likelihood of which depends on the realized values of these variables. Since $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ cannot be known in advance, the fulfillment of inequality (16) is uncertain. Larger parameter values have a lower likelihood of yielding improvements, although such improvements can be substantial, whereas smaller parameter values increase the likelihood of improvement, albeit potentially marginal or negligible, compared with the naïve forecast. Therefore, this boundary criterion highlights a critical trade-off between the likelihood of achieving improvements and the potential magnitude of these improvements.
In addition to using $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ as the estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, a more conservative approach can be used to generate the estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$:
@@
    Within each window, the directional accuracy is calculated and then translated into the corresponding value of $\theta$ via Eq. (9). As the window moves, a collection of $\theta$ values is generated, denoted as $\Theta_{\text{set}}$, whose bounds can be regarded as the empirical range of $\theta_{\text{out-of-sample}}^{*}$.
@@
    Similarly, a sliding window of the same length is applied to calculate a set of $\bar{\epsilon}$ values from the in-sample set, denoted as $\boldsymbol{\bar{\epsilon}}_{\text{set}}$, providing an empirical value range of $\bar{\epsilon}_{\text{out-of-sample}}$.
@@ 3 Simulation
This section presents a simulation study using four synthetic time series to evaluate whether the adjusted naïve forecast outperforms the original naïve forecast when movement predictions with different directional accuracy levels are provided. The performance of the proposed method under each directional accuracy level was evaluated based on the out-of-sample set of each time series.
@@ 3.1 Datasets and Setup
The dataset includes four synthetic time series, each of which represents a symmetric random walk with a certain type of variance, as summarized in Table 1. Each time series, consisting of 2,500 steps, was generated using a fixed random seed (seed = 1) in the Python environment, as shown in Figure 2. To maintain all values within the positive range, an initial positive offset of 10000 units was applied to each series. The last 500 data points were selected as the out-of-sample set, and the remaining data points were used as the in-sample set. In addition, Figure 3 shows the first-order differences of each time series to visually demonstrate the properties of the increments, and Figure 4 shows the histograms of the first-order differences for all time series, highlighting the symmetric nature of the time series.
Table 1: Variance of synthetic random walks
@@
For each time series, the adjusted naïve forecast on the out-of-sample set was generated via Algorithm 1. The performance of each adjusted naïve forecast result was evaluated via four metrics: the RMSE, MAE, MAPE, and sMAPE. The RMSE, derived from the square root of the MSE, emphasizes larger errors, which makes it suitable for identifying significant deviations in forecasts. The MAE, on the other hand, provides an interpretable measure of the average error without overpenalizing outliers. The MAPE and sMAPE normalize the errors relative to the actual values, enabling comparisons between datasets with different scales. Together, these metrics capture both absolute and relative error characteristics, ensuring a robust and balanced evaluation. The formulas for these metrics are as follows:
@@
To statistically test the performance of the adjusted naïve forecast, for each synthetic time series at a specified directional accuracy level, the process to generate the adjusted naïve forecast was repeated 100 times. Each repetition uses a unique movement prediction sequence at the same directional accuracy level, which is generated by randomly replacing a subset of actual movements in the out-of-sample set with incorrect movements. The directional accuracy level varies in two ways: initially, it changes from 0.50 to 0.56 in increments of 0.01, and then it changes from 0.55 to 1.00 in larger increments of 0.05. This approach allows for finer granularity in the lower accuracy range, where small changes in the directional accuracy can have a significant effect on the performance of the method, while using larger steps at higher accuracy levels balances the computational efficiency. Consequently, for each synthetic time series, 100 independent values were calculated for each metric at each accuracy level.
@@ 3.2 Results and Analysis
Because the directional accuracy for the out-of-sample set is predefined in each simulation process, the value of $\theta_{\text{out-of-sample}}^{*}$ can be derived directly via Eq. (14), and only the value of $\bar{\epsilon}_{\text{out-of-sample}}$ needs to be estimated. Figure 5 illustrates the distribution of the values in the $\boldsymbol{\bar{\epsilon}}_{\text{set}}$, generated based on the sliding windows across the first-order differences of the in-sample set. The vertical reference line in Figure 5 represents the actual value of $\bar{\epsilon}_{\text{out-of-sample}}$. In the four cases, the minimum value of $\boldsymbol{\bar{\epsilon}}_{\text{set}}$ is generally an underestimate of $\bar{\epsilon}_{\text{out-of-sample}}$, except for the second case, where the synthetic time series is a random walk with a decreasing trend of linear variance. In general, the estimated value remains close to the true value. Under the assumption that the directional accuracy is consistent across the in-sample and out-of-sample sets, this estimation ensures that inequality (16) is satisfied.
@@
Furthermore, comparisons of the two types of naïve forecasts based on the four evaluation metrics at low accuracy levels are illustrated in Figures 6, 7, 8, and 9. Each box plot in each figure represents the distribution of 100 independent values for one metric at a specific directional accuracy level, with the metric value of the original naïve forecast shown as a reference line. According to Figure 6, at low accuracy levels, such as 0.51, the RMSE of the adjusted naïve forecast is comparable to or slightly better than that of the original naïve forecast, as shown by the proximity of the median of each box plot to the reference line. As the accuracy increased beyond 0.52, most of the RMSE values fell below the baseline.
@@
Because the adjusted naïve forecast consistently outperforms the original naïve forecast at low accuracy levels in terms of the MAE, MAPE, and sMAPE, the Wilcoxon signed rank test was conducted to specifically evaluate whether the median differences between the RMSE values of the adjusted naïve forecast and those of the original naïve forecast significantly deviated from zero for accuracy levels ranging from 0.50–0.56 (Li and Johnson,, 2014). As summarized in Table 2, at a directional accuracy level of 0.51, the adjusted naïve forecast shows mixed results with respect to statistical significance. Although significant improvements were observed over the original naïve forecast for the Type 4 series (randomly varying variance, p = 0.002), the results for the Type 1 (constant variance, p = 0.915), Type 2 (linear variance trend, p = 0.781), and Type 3 (cyclic variance, p = 0.812) series were not statistically significant, indicating that the adjusted naïve forecast under an accuracy level of 0.51 may not consistently produce meaningful improvements across all types of time series. As the accuracy increased beyond 0.51, significant reductions in the RMSE (p < 0.001) were observed for all types of time series, thereby demonstrating the reliability and effectiveness of the adjusted naïve forecast at higher accuracy levels.
@@
According to Figures 10, 11, 12, and 13, at higher accuracy levels ranging from 0.55–1.00, the adjusted naïve forecast consistently outperforms the original naïve forecast in terms of the four metrics. In particular, the improvement over the original naïve forecast becomes substantial as the directional accuracy improves. This finding demonstrates that the proposed method can effectively translate highly accurate movement predictions into improved point forecasts in the context of forecasting random walk-like time series. However, the improvement over the original naïve forecast has a limit. The proposed method allows a random walk-like series to be predicted to a certain extent, but the predicted value cannot be infinitely close to the actual value.
@@
In general, when the directional accuracy was 0.5, the adjusted naïve forecast did not provide an improvement over the original naïve forecast. As the directional accuracy increases, the performance of the adjusted naïve forecast improves. The simulation results also demonstrate that the effectiveness of the proposed method is unaffected by the variance type of the synthetic time series, as evidenced by the consistent performance of the adjusted naïve forecast across all synthetic series. This result is consistent with the theoretical analysis of the proposed method, which highlights its robustness and potential applicability to a wide range of time series that exhibit symmetric random walk characteristics, including those with homoscedasticity as well as high-volatility scenarios.